The air temperature is 20 oC. A stationary observer sees a police car receding into the distance with a speed of 30 m/s with it’s siren blaring. If the police siren is designed to produce a sound of frequency 900 Hz, calculate the frequency of the sound heard by the stationary observer.


1
Expert's answer
2021-01-29T11:58:29-0500
"f'=f\frac{v-v_o}{v+v_s}\\f'=900\frac{330-30}{330+0}=818\ Hz"